Stat. class. How do you consruct a pie chart when only percentages are given? I don't have frequencies to compute degrees. Components of Earth's crust:
O2 45.6%
Silicon 27.3
Alum. 8.4
Fe 6.2
Ca 4.7
Other 7.8
I have a TI-83 plus but I don't know if it'll help, or how to use it yet.

Answer by Edwin McCravy(20086)

Element    Percent     Percent of 360° = Angle 
O2          45.6%          .456 × 360° = 164.16° 
Silicon     27.3           .273 × 360° =  98.28°
Alum.        8.4           .084 × 360° =  29.88°
Fe           6.2           .062 × 360° =  22.32°
Ca           4.7           .047 × 360° =  16.92°
Other        7.8           .078 × 360° =  28.08°

Draw the pizza, using a compass for the circle and a protractor
to measure the central angles.  You might also put the 
percentages in the pieces of pie even though I didn't here:

             

Edwin